Add: x2 – y 2 – 1, y2 – 1 – x 2 , 1 – x 2 – y 2 Add: x2 – y 2 – 1, y2 – 1 – x 2 , 1 – x 2 – y 2

Add: 
 x2 – y 2 – 1, y2 – 1 – x 2 , 1 – x 2 – y 2
Add: 
 x2 – y 2 – 1, y2 – 1 – x 2 , 1 – x 2 – y 2

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

When term have the same algebraic factors, they are like terms. Then, we have to add the like terms
= x^2–y^2-1+ (y^2-1-x^2) + (1-x^2-y^2)
= x^2–y^2-1+ y^2-1-x^2 + 1-x^2-y^2
= x^2-x^2-x^2-y^2+y^2-y^2-1-1+1
=x^2 (1-1-1)+ y^2(-1 + 1 - 1) + (-1 -1 + 1)
= x^2(1-2)+ y^2(-2 +1) + (-2 + 1)
= x^2(-1) + y2(-1) + (-1)
= -x2–y2-1

Thanks